OpenStudy (anonymous):

How do you solve When the lesser of two consecutive integers is added to three times the greater the result is 43

Directrix (directrix):

Let x and x + 1 be the consecutive integers. x is the less one and x+ 1 is the greater one. Three times the greater is written in symbols as 3( x + 1 ) the lesser of two consecutive integers is added to three times the greater the result is 43 Translating the words gives the following: x + 3 ( x + 1 ) = 43 x + 3x + 3 = 43 4x + 3 + 43 4x = 40 x = 10 Lesser integer is 10; the greater integer is 11.